The difference between bases on a square baseball field is 90 feet. To the nearest tenth, what is the straight-line distance, in feet, from first base to third base?


Answer by checkley79(3341)

A^2+B^2=C^2
90^2+90^2=C^2
8,100+8,100=C^2
16,200=C^2
C=SQRT16,200
C=127.3 FEET ANS.